Glyphosate, the active ingredient in Roundup herbicide, degrades by a first-order process having ahalf-life of 32 days. After 96 days,

(A)

the fraction of glyphosate remaining cannot be

determined without knowing the initial

concentration of glyphosate.

(B)

about 88% of the glyphosate remains.

(C)

about 67% of the glyphosate is degraded.

(D)

about 12% of the glyphosate remains.

(E)

about 25% of the glyphosate remains.

Please explain !
Answer #1

96 days = 96/32 = 3 half lives

After N half lives, the fraction is:

N/No = (1/2)^n

N/No = (1/2)^3

N/No = 1/8

So,

N = No/8

So, if No was 100 %

N = 100/8 = 12.5 %

That is 12.5 % remains

Answer: D